OpenStudy (anonymous):

hey i could use a hand on this problem cant wrap my head around the process. Use the points (0, 4) and (-2, 3) to find the slope of the line and give the equation of this line

OpenStudy (anonymous):

slope= (3-4)/(-2-0) =1/2 equation of line y-4=1/2(x-0) y-4=x/2 2y-4=x

OpenStudy (anonymous):

thnx your awsome